Today:
- installed rad and soils in corn at rel, plugged into rel.a1, USB7. There are a few bits of nearby corn plants that are higher than the radiometer, even mounted as high as I dared on the 10' Rohn tower section.
 - took a long break with a telecon and lunch
 - added NR01 in soybeans at lconv
 - installed rad and soils in grass at uconv, plugged into a2, USB7. THIS COMPLETES SENSOR INSTALLATION!
 - Did some ATV moving of PI equipment
 - Cleaned our sites of extraneous stuff
 - Starting sensor status, seeing a bunch of stuff
- uconvm, lconv1, lconvt needed ddn/dup to appear in cockpit
 - still 10m csat3 on uconvm is not responding
 - lots of cockpit scales need to be set (easy addition to .xml)
 - 10m csat3 on lconvm is reporting bad data
 - 6m init csat3 in bad cycle mode (needs a head swap – our last spare (we had one other bad already))
 - ...and I'm sure there's more 
 
 - A visit to storage with a few things
 
Tonight:
- make obvious .xml edits, push to git
 
Tomorrow:
- Kate and I are going to visit UI Atmos. Sci. in the morning to visit Steve Nesbit and Francina
 - ansible push from git (I'll have to lookup the command)
 - cable clean-up
 - address the inevitable sensor data issues
 - clean up the base and start working out of it